A meter rule is hung by two light strip tied at the 15cm, and 75 cm mark respectively to a
ceiling. If the weight of the ruler is 2N. (i) calculate the tension in the two string.
1
Expert's answer
2021-09-02T10:42:20-0400

Draw the diagram:



According to equilibrium of forces:


"T_A+T_B-W=0,\\\space\\
T_A+T_B=W=2\text N."

The total tension in the two strings is 2N.


Find the torque around point A:


"\frac{15}{2}·2·(15/100)-\frac{85}{2}·2·(85/100)+60·T_B=0,\\\space\\
T_B=1.17\text N.\\
T_A=2-T_B=0.83\text N."